OpenStudy (anonymous):

2x+3y=0 8x+5=5y solve with elimination

OpenStudy (anonymous):

\[\left( -\frac{ 15 }{ 34 }, \frac{ 5 }{ 17}\right)\]

OpenStudy (anonymous):

is what i got